# Expected Gamma-Ray Burst Detection Rates and Redshift Distributions for the BlackCAT CubeSat Mission

### v1

@@ Abstract
We report the results of an extensive set of simulations exploring the sensitivity of the BlackCAT CubeSat to long-duration gamma-ray bursts (GRBs). BlackCAT is a NASA APRA-funded CubeSat mission for the detection and real-time sub-arcminute localization of high-redshift () GRBs. Thanks to their luminous and long-lived afterglow emissions, GRBs are uniquely valuable probes of high-redshift star-forming galaxies and the intergalactic medium. In addition, each detected GRB with a known redshift serves to localize a region of high-redshift star formation in three dimensions, enabling deep follow-on searches for host galaxies and associated local and large-scale structure. We explore two distinct models for the GRB redshift distribution and luminosity function, both consistent with Swift observations. We find that, for either model, BlackCAT is expected to detect a mean of 42 bursts per year on-orbit, with 6.7% to 10% of these at . BlackCAT bursts will be localized to precision and reported to the community within seconds. Due to the mission orbit and pointing scheme, bursts will be located in the night sky and well-placed for deep multiwavelength follow-up observations. BlackCAT is on schedule to achieve launch readiness in 2025.
@@ 1 Introduction
Successful launch and commissioning of JWST (Gardner et al., 2023) has provided the global astronomical community with a revolutionary new set of capabilities for studying the high-redshift universe. As a result, from the moment of the mission’s first data release, analyses have proliferated identifying (Finkelstein et al., 2023; Carnall et al., 2023; Hainline et al., 2023; McLeod et al., 2023; Austin et al., 2023; Casey et al., 2023) and spectroscopically confirming (Fujimoto et al., 2023; Arrabal Haro et al., 2023; Curtis-Lake et al., 2023) a rich harvest of massive star-forming galaxies from redshifts . These indications of vigorous early star-forming activity have been bolstered by JWST identification of massive and merging galaxies at slightly lower redshifts, (Labbé et al., 2023; Boyett et al., 2023).
While the quantitative picture of this dawn of star formation remains under construction, there is already widespread agreement that star formation was underway early, energetically, and extensively well before the end of what had previously been termed the “cosmic dark ages” (Bouwens et al., 2022; Wang et al., 2023; Leung et al., 2023; Austin et al., 2023; Mascia et al., 2023; Donnan et al., 2023). In particular, there is no sign yet of a sharp cutoff to star formation in the early universe, neither at the end of reionization at nor at the highest redshifts probed to date by JWST programs.
@@
GRBs act as reliable tracers of star formation at (Greiner et al., 2015; Schulze et al., 2015; Sears et al., 2023). Since the Neil Gehrels Swift Observatory (Gehrels et al., 2004) has discovered GRBs with spectroscopic redshifts out to (Tanvir et al., 2009; Salvaterra et al., 2009; Tanvir et al., 2018) and photometric redshifts out to (Cucchiara et al., 2011), the existing burst sample already extends into the era from which JWST is collecting statistical samples of luminous star-forming galaxies.
@@ 3.1 GRB Population Simulations
The case for weak evolution rests on the metallicity argument for differential evolution at low redshift, along with observations of GRB host galaxies (Greiner et al., 2015; Schulze et al., 2015; Sears et al., 2023), which are consistent with these host galaxies being drawn (as weighted by galaxies’ star-formation rates) from the star-forming galaxy population.

### v2

@@ Abstract
We report the results of an extensive set of simulations exploring the sensitivity of the BlackCAT CubeSat to long-duration gamma-ray bursts (GRBs). BlackCAT is a NASA APRA-funded CubeSat mission for the detection and real-time sub-arcminute localization of high-redshift () GRBs. Thanks to their luminous and long-lived afterglow emissions, GRBs are uniquely valuable probes of high-redshift star-forming galaxies and the intergalactic medium. In addition, each detected GRB with a known redshift serves to localize a region of high-redshift star formation in three dimensions, enabling deep follow-on searches for host galaxies and associated local and large-scale structures. We explore two distinct models for the GRB redshift distribution and luminosity function, both consistent with Swift observations. We find that, for either model, BlackCAT is expected to detect a mean of 42 bursts per year on-orbit, with 6.7% to 10% of these at . BlackCAT bursts will be localized to precision and reported to the community within seconds. Due to the mission orbit and pointing scheme, bursts will be located in the night sky and well-placed for deep multiwavelength follow-up observations. BlackCAT is on schedule to achieve launch readiness in 2025.
@@ 1 Introduction
Successful launch and commissioning of JWST (Gardner et al., 2023) has provided the global astronomical community with a revolutionary new set of capabilities for studying the high-redshift universe. As a result, from the moment of the mission’s first data release, analyses have proliferated identifying (Finkelstein et al., 2023; Carnall et al., 2023; Hainline et al., 2024; McLeod et al., 2024; Austin et al., 2023; Casey et al., 2024) and spectroscopically confirming (Fujimoto et al., 2023; Arrabal Haro et al., 2023; Curtis-Lake et al., 2023) a rich harvest of massive star-forming galaxies from redshifts . These indications of vigorous early star-forming activity have been bolstered by JWST identification of massive and merging galaxies at slightly lower redshifts, (Labbé et al., 2023; Boyett et al., 2024).
While the quantitative picture of this dawn of star formation remains under construction, there is already widespread agreement that star formation was underway early, energetically, and extensively well before the end of what had previously been termed the “cosmic dark ages” (Bouwens et al., 2023; Wang et al., 2023; Leung et al., 2023; Austin et al., 2023; Mascia et al., 2024; Donnan et al., 2023). In particular, there is no sign yet of a sharp cutoff to star formation in the early universe, neither at the end of reionization at nor at the highest redshifts probed to date by JWST programs.
@@
GRBs act as reliable tracers of star formation at (Greiner et al., 2015; Schulze et al., 2015; Sears et al., 2024). Since the Neil Gehrels Swift Observatory (Gehrels et al., 2004) has discovered GRBs with spectroscopic redshifts out to (Tanvir et al., 2009; Salvaterra et al., 2009; Tanvir et al., 2018) and photometric redshifts out to (Cucchiara et al., 2011), the existing burst sample already extends into the era from which JWST is collecting statistical samples of luminous star-forming galaxies.
@@ 3.1 GRB Population Simulations
The case for weak evolution rests on the metallicity argument for differential evolution at low redshift, along with observations of GRB host galaxies (Greiner et al., 2015; Schulze et al., 2015; Sears et al., 2024), which are consistent with these host galaxies being drawn (as weighted by galaxies’ star-formation rates) from the star-forming galaxy population.